### v2

# ControlDreamer: Stylized 3D Generation with Multi-View ControlNet

Abstract
Recent advancements in text-to-3D generation have significantly contributed to the automation and democratization of 3D content creation. Building upon these developments, we aim to address the limitations of current methods in generating 3D models with creative geometry and styles. We introduce multi-view ControlNet, a novel depth-aware multi-view diffusion model trained on generated datasets from a carefully curated $100K$ text corpus. Our multi-view ControlNet is then integrated into our two-stage pipeline, ControlDreamer, enabling text-guided generation of stylized 3D models. Additionally, we present a comprehensive benchmark for 3D style editing, encompassing a broad range of subjects, including objects, animals, and characters, to further facilitate diverse 3D generation. Our comparative analysis reveals that this new pipeline outperforms existing text-to-3D methods as evidenced by qualitative comparisons and CLIP score metrics.
Keywords:
text-to-3D, two-stage pipeline, multi-view diffusion model, ControlNet
**footnotetext: These authors contributed equally to this work\$$\dagger$\$$\dagger$\$footnotetext: Corresponding author
1 Introduction
3D content creation is recently gathering attention, particularly in the realms of virtual reality and game development. This evolving landscape has been profoundly impacted by the recent advancement of 2D lifting methods. Dreamfusion’s [29] introduction of score distillation sampling (SDS) has revolutionized and democratized the field, facilitating the generation of 3D models by leveraging large-scale text-to-image diffusion models [35, 33, 1]. This breakthrough has enabled more intuitive and imaginative 3D content creation from user’s textual descriptions. Furthering this evolution, various works show high-resolution text-to-3D generation, employing multiple 3D representations [21, 4, 5] or variants of score distillation sampling [42, 11].
However, current 2D lifting methods face the Janus Problem of generating multiple faces from different angles due to a lack of 3D knowledge. MVDream [39] addresses this by adapting Stable Diffusion’s [33] 2D self-attention to 3D and jointly training with multi-view images from the Objaverse [8] and LAION datasets [36]. For instance, prompts like ‘Captain America’ tend to yield 3D model of a shield rather than the character, likely due to dataset biases and the mode-seeking nature of score distillation sampling [29]. This leads to the generation of mode geometries learned from the dataset, which limits the creation of customized 3D models, such as a bulky Captain America (Fig. 1).
Figure 1: Comparing text-to-3D pipeline of our ControlDreamer and MVDream. The right side showcases MVDream’s output, which demonstrates a vulnerability to geometry biases learned during pre-training, often resulting in (a) incomplete models like only generating Captain America’s shield or a lioness’s head or (b) stereotypical geometries related to the prompts. In contrast, the left side displays how ControlDreamer, through its two-stage pipeline, effectively overcomes geometry biases, enabling unique combinations of geometry and style in 3D models, even for counterfactual generation.
These challenges motivated us to study a two-stage approach given two text prompts. Initially, we generate a neural radiance field (NeRF) [23, 24] from a geometry prompt to sculpt geometric structures, followed by converting it to a mesh using DMTet algorithm [38] and refining both geometry and texture using a style prompt. In the second stage, we found that pre-trained MVDream model encounters difficulties in aligning textures with the initially generated geometrical structures, resulting in noticeable artifacts as represented in Fig. 2 (b). To address this, we present MV-ControlNet, a depth-aware multi-view diffusion model. We first generate a paired dataset of text and multi-view images using GPT-4 filtered texts and MVDream, followed by training a depth conditioner for the frozen MVDream, according to the principle of ControlNet [43]. Obtaining MV-ControlNet is cost-effective, requiring only 1.5 days on dataset generation and 1.5 days on training, on four A40 GPUs (Note, MVDream required 3 days on 32 A100 GPUs for model training).
In evaluations conducted on our new benchmark, which consists of pairs of geometry and style prompts, ControlDreamer demonstrates superior performance in generating styles on 3D models compared to existing two-stage pipelines. By using CLIP scores and human assessments, we find that our depth-aware MV-ControlNet outperforms normals and edges-aware variants in rendering detailed textures and geometries. These findings confirm ControlDreamer’s ability to produce high-quality 3D models that closely match the provided textual descriptions. For more detailed results and comparisons, please visit our project page.
Our contributions can be summarized as follows:
  • We introduce ControlDreamer, a two-stage pipeline for 3D model generation, combining NeRF creation from geometry prompts and mesh refinement with style prompts.
  • The training of MV-ControlNet is more practical and resource-efficient, offering improvements in time and resource usage compared to prior models.
  • Our MV-ControlNet adeptly aligns diverse geometries and styles, elevating the generation of view-consistent geometry and styles from various texts.
2 Related Works
2.1 Adding Conditions to Text-to-Image Models
Diffusion models like Stable Diffusion [33], Imagen [35], and eDiff-I [1], trained on billions of image-text pairs, have shown remarkable capabilities in generating high-quality images from text. ControlNet [43] introduces a method for adding conditions to pre-trained text-to-image models by using a trainable copy alongside the original frozen model for specific tasks like depth-to-image. Instruct-Pix2Pix [2] uses GPT-3 [3] and Stable Diffusion to create a training dataset for an image editing model that follows diverse editing instructions.
Figure 2: Results from generating various stylized 3D models using style prompts, based on a source NeRF created from a geometry prompt. Our ControlDreamer (a) aligns various styles on the source geometry, while the baseline (b) suffers artifacts due to misalignment of style and geometry.
2.2 Text-to-3D Generation
Training a generative 3D model using large-scale 3D datasets is highly challenging due to the difficulty and expense of collecting 3D data compared to images. DreamFusion [29] introduces the score distillation sampling (SDS), enabling the optimization of an implicit 3D model [23, 24] without any 3D dataset by utilizing a pre-trained 2D diffusion model as a prior. Despite these advantages, DreamFusion has the drawback of generating scenes very slowly and producing lower-quality results. To overcome these limitations, several methods are proposed involving primary optimization for NeRF and secondary optimization for the extracted mesh [21, 5], along with improvements for SDS [42, 15].
Table 1: Previous text-to-3D generation scennarios. Our method is the first two-stage pipeline designed for multi-view consistent 3D generation. Notably, it offers convenience by eliminating the need for annotations like bounding boxes and masks.
Method 3D Rep. Multi-Stage Annot.-Free Multi-View
Vox-E [37] Voxel
PG3D [6] NeRF
TEXTure [32] Texture
Text2tex [4] Texture
ProlificDreamer [42] NeRF
Fantasia3D [5] DMTet
Magic3D [21] NeRF&\&&DMTet
MVDream [39] NeRF
Ours NeRF&\&&DMTet
2.3 Multi-view Diffusion Nodel
In contrast to prior 2D lifting-based methodologies, recent advancements have introduced multi-view aware training for text-to-image diffusion models. MV-diffusion [40] facilitates the generation of multi-view panoramic images and depth-to-image conversions using given text prompts. This approach introduces a novel attention module to account for pixel-to-pixel correspondences, ensuring coherence across different views. MVDream [39] integrates a 3D self-attention module and fine-tunes the diffusion model through joint training on large-scale 2D and 3D datasets. As a result, MVDream avoids the Janus Problem and Content Drifting Problem, consistently producing 3D models that align accurately with the provided texts.
2.4 Text-Guided 3D Editing
In text-to-3D generation, various innovative methods have been proposed. Vox-E [37] edits voxel grids with SDS, maintaining geometry and appearance. Instruct-NeRF2NeRF [11] iteratively updates source NeRFs using Instruct-Pix2Pix [2]. Text2tex [4] assumes a perfect mesh, applying textures without mesh refinement flexibility. Magic3D [21] offers a coarse-to-fine pipeline for high-resolution 3D models, starting with NeRF and converting to deformable meshes. Fantasia3D [5] divides its process into geometry and appearance stages, beginning with a 3D ellipsoid and refining with DMTet. Progressive3D [6] decomposes complex prompts for localized editing, focusing on user-specified regions.
In Table 1, we outline a key gap in existing research: a lack of methods combining multi-stage, annotation-free, and multi-view capabilities for 3D generation. We found that current models, especially in DMTet texture refinement, work best when geometry and style prompts match. Furthermore, methods using 2D models struggle with multi-view consistency. Addressing this, we integrated multi-view features from MVDream in our ControlDreamer, enabling high-quality, differentiated prompt-based 3D generation.
3 Method
Our ControlDreamer pipeline involves a two-stage process: first, training a coarse-grained NeRF with a geometry prompt and converting it into a textured mesh using DMTet, and then creating a fine-grained stylized 3D model with a style prompt. We illustrate our pipeline in Fig. 3 and provide detailed explanations in Sec. 3.1. Then, we emphasize a depth-aware, multi-view approach in the second stage and elaborate on the Multi-View (MV) ControlNet’s training scheme in Sec. 3.2.
Figure 3: Illustration of ControlDreamer. (Left) Using a geometry prompt and MVDream, we create a multi-view consistent NeRF, with MVDream’s 3D self-attention ensuring consistency. (Right) The NeRF is transformed into a mesh via DMTet and undergoes style generation using our MV-ControlNet, which combines a trainable copy (red) and zero convolution (white). MV-ControlNet is designed to understand geometry from the multi-view depths, enabling effective style generation.
3.1 ControlDreamer: Two-Stage 3D Generation
3.1.1 Geometry Stage
Here, we generate a NeRF from a given geometry prompt. The style of this NeRF will subsequently be modified in the second stage. We optimize the NeRF by distilling probability density [26] from the pre-trained MVDream [39]. This optimization is achieved through score distillation sampling (SDS) [29], where the diffusion model directly computes the loss gradient as shown in the following equation:
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}\triangleq\mathbb{E}_{t,\epsilon}\left[w(t)\left(\hat{\epsilon}_{\phi}(z_{t};y_{geo},c,t)-\epsilon\right)\frac{\partial z}{\partial x}\frac{\partial x}{\partial\theta}\right],$ (1)
where $t\sim U(t_{\text{min}},t_{\text{max}})$, $w(t)$ denotes weighting function, $\hat{\epsilon}_{\phi}(z_{t};y,t)$ represents a guided epsilon prediction [14], $y_{geo}$ is the geometry text prompt, $c$ represents camera parameter, $z$ refers to VAE latents [18, 33] of rendered multi-view images $x$, and $\theta$ is the parameter of 3D model. Note that there is no backpropagation through the diffusion model.
To leverage the multi-view consistency of MVDream, we render four views from randomly selected camera parameters $c$ in each iteration of optimization. These parameters are uniformly sampled from a range of view angles, all maintaining a consistent elevation to ensure a diverse yet coherent range of perspectives. In further detail, we incorporate lighting strategies such as point lighting [29], orientation loss [29], and soft shading [21]. Additionally, we enhance the background with random color augmentation and employ a fixed negative prompt to avoid low-quality 3D models. The timestep range $[t_{\text{min}},t_{\text{max}}]$ is linearly annealed during the optimization process. To prevent color saturation, we apply the $x_{0}$-reconstruction loss, which is the epsilon reconstruction loss weighted by the signal-to-noise ratio (SNR) [17], thus $w(t)=\text{SNR}(t)$. The text prompt $y_{geo}$ is embedded using the CLIP ViT-L text encoder [30].
3.1.2 Style Stage
Our second stage is motivated by the coarse-to-fine optimization of [21], which is utilized for generating high-resolution meshes. In their process, the textured mesh is optimized with DMTet algorithm [38], which uses a deformable tetrahedral grid and employs a differentiable rasterizer [19, 25] for efficient rendering. However, unlike the goal of [21], which is centered on high-resolution mesh refinement, our approach leverages the DMTet algorithm for generating style on the previously generated 3D model.
In each iteration of optimization, we employ pre-trained MiDAS [31] to obtain predicted depth maps, leveraging the identical depth estimator that is used for training our MV-ControlNet in Sec. 3.2. Subsequently, we optimize mesh with the gradient computed by MV-ControlNet as shown in the following equation:
$\nabla_{\theta}\mathcal{L}_{\text{SDS}-\text{C}}\triangleq\mathbb{E}_{t,\epsilon}\left[w(t)\left(\epsilon_{\phi}(z_{t};\mathcal{D}(x),y_{style},c,t)-\epsilon\right)\frac{\partial z}{\partial x}\frac{\partial x}{\partial\theta}\right],$ (2)
where $y_{style}$ is the style text prompt, which is differently used with $y_{geo}$. $\mathcal{D}$ is our MV-ControlNet, which we later explain in Sec. 3.2. This gradient is coupled with gradients of regularizations to meticulously refine both geometry and texture. In contrast to recent methods [21, 5, 42] that necessitate narrow $[t_{\text{min}},t_{\text{max}}]$ during the refinement phase, our ControlNet allows a large range of $[t_{\text{min}},t_{\text{max}}]$ with scheduled timestep annealing. This ensures that depth information remains intact, safeguarding original geometry throughout the style optimization, even with significant variations in timesteps.
Similar to the iterative dataset update concept introduced in [11], we discovered that iteratively updating mesh with MiDAS-guided estimated depth maps effectively edits the geometry and texture during a single-stage training of the given source 3D models toward the desired target 3D models.
3.1.3 Regularizations
We use the following two regularizations along with score distillation sampling. First, we use normal map consistency loss [21], which ensures the mesh’s normal map remains consistent. Second, we use Laplacian smoothing regularization [42], which eliminates pixelated artifacts in DMTet. These allow us to generate photorealistic 3D models.
3.1.4 Annealing diffusion timesteps
During optimization, we linearly anneal the timestep range of $t$. Utilizing MV-ControlNet, we have discovered that textures can be effectively modified while preserving the initial geometry, even with significant variations in text descriptions. Furthermore, this effectiveness persists in scenarios where $t$ is sampled from $t\sim U(0.02,0.98)$. We continued sampling $t$ from this range for the initial $4/5$ of the total iterations, akin to the geometry stage. Subsequently, for the remainder of the iterations, $t$ is sampled from $t\sim U(0.02,0.5)$. MV-ControlNet allows us to deviate from traditional methods that rely on smaller timesteps for mesh refinement.
Figure 4: Additional qualitative results. Our generation process involves first generating a coarse-grained geometry, and then generating a fine-grained stylized 3D model with a style prompt.
3.1.5 Editing the geometry and texture
ProlificDreamer [42] comprises two refinement stages: geometric refinement followed by texture refinement. Furthermore, each training iteration requires not only the computation of the SDS loss but also the fine-tuning of the diffusion model, resulting in a computationally expensive generation process.
3.2 Multi-View ControlNet
As elaborated in the previous section, our pipeline generates style in the second stage by refining the DMTet, which is converted from the NeRF. We note the limitations of MVDream in refining DMTet, stemming from a lack of understanding of the geometry. To address this, we introduce the Multi-View (MV) ControlNet, crucial for generating fine-grained style while comprehending the coarse-grained mesh’s geometry.
3.2.1 Generating Dataset
To create a depth-aware multi-view model, we use the curated 100K high-quality texts from OpenShape [22], filtered with GPT-4 [27] for quality assurance. These texts are first refined using BLIP [20] and Azure cognition services for initial 2D image captions, and then enriched by GPT-4, filtering out uninformative content. We utilize MVDream-generated multi-view images as training data, matching MVDream’s camera parameters to maintain its prior knowledge. This involves generating four orthogonal views for each text, with uniformly distributed azimuth angles and elevation. Specifically, we randomly sample the camera parameters in a range of [0.9, 1.1] for distance, [15, 60] for fov, and [0, 30] for elevation. Then, we generate multi-view depth maps from the generated images using MiDAS [31]. The process, completed in 1.5 days on four A40 GPUs, utilizes texts with explicit ‘3D’ mentions or appended them with a ‘3D asset’ postfix.
3.2.2 Training MV-ControlNet
MVDream [39] expands the text-to-image capabilities of Stable Diffusion [33] to enable text-to-multi-view generation. The model processes inputs of noisy latents $z_{t}\in\mathbb{R}^{F\times H\times W\times C}$, camera parameters $c\in\mathbb{R}^{F\times 16}$, and text prompts. MVDream differentiates itself by utilizing a 3D self-attention mechanism, an extension of 2D self-attention, to model multi-view consistency and capture the interrelations among pixels across $F$ views. Note that additional operations such as cross-attentions and convolutions operate on a frame-wise basis.
Our goal is to train a depth encoder that incorporates multi-view depth conditions into MVDream, leveraging the model’s multi-view consistency and generation quality. Following the ControlNet [43] approach, which involves training an additional encoder atop a frozen diffusion model, we introduce multi-view depths as additional inputs to the pre-trained MVDream. Specifically, a trainable copy of MVDream’s U-Net [34] encoder, including 3D attentions, is tasked with encoding depth and extracting multi-resolution features, which are then integrated into the U-Net decoder’s skip connections using zero-initialized convolutions. We train this encoder with the epsilon prediction objectives [13], formalized as follows:
$\mathcal{L}=\mathbb{E}_{z\sim\mathcal{E}(x),t,\epsilon\sim\mathcal{N}(0,1)}\left[\left\|\epsilon-\epsilon_{\phi}(z_{t};\mathcal{D}(x),y,c,t)\right\|^{2}\right],$ (3)
where $\mathcal{E}$ denotes the VAE encoder of Stable Diffusion [33], $\mathcal{D}$ is the trainable depth encoder, $\epsilon_{\phi}$ is the frozen MVDream’s U-Net, $y$ is the text prompt, and $c$ is the camera parameter.
Figure 5: We compare the depth-aware MV-ControlNet with [12] approach using MVDream, as well as MV-ControlNet variants trained under the edges and normals conditions. The left side shows source images along with their conditions. The depth-conditioned multi-view images consistently display the most visually appealing results.
Figure 6: The left column displays an array of initial 3D models. From each initials, multiple styles are applied to create 3D models. Utilizing a two-stage pipeline, this process generates customized 3D models with a variety of geometries, including samurai, Mr. Bean, stone bust, and a Messi figure, and styles including Iron Man, Captain America, and Superman. Our ControlDreamer demonstrates visually appealing results for diverse, and challenging concepts.
4 Experiments
In our experiments, we carry out tests using user-customized prompts, distinct from the $100K$ text corpus used for training our MV-ControlNet. For both qualitative and quantitative assessments, we have created a benchmark dataset comprising a total of 30 pairs of geometry and style prompts specifically for 3D editing experiments. These pairs span across five distinct categories: Animals, Characters, Foods, General, and Objects. The main qualitative samples of our pipeline are presented in Fig. 6.
4.1 Qualitative Comparison of Multi-View Image Generation
To evaluate our MV-ControlNet’s depth-aware multi-view image generation capabilities, we conduct comparisons with several baselines as shown in Fig. 5. The baselines include MVDream using the prompt-to-prompt (P2P) [12] method and MV-ControlNet variants trained under the canny edge and normal conditions. P2P is an image editing method that aligns source and target images’ geometries by injecting attention maps into diffusion models. We use injection ratios set at 0.4 for cross-attention and 0.8 for self-attention, ensuring strong geometric alignment. However, the applicability of P2P is somewhat constrained due to the need for equal token lengths in both source and target texts, often resulting in less aligned text-guided editing outcomes. Conversely, as the proposed MV-ControlNet can ensure source geometry with depth conditions, it is not restricted by the length of the text, enabling the user’s application of more creative and diverse prompts in the style stage. Additional comparative results are displayed in Fig. 15 in the Appendix.
Regarding MV-ControlNet variants, we observe that canny edge maps result in unrealistic and 3D-inconsistent images, whereas normal maps provide 3D consistency but show degraded text alignment. Conversely, using depth conditions is the most effective, leading to visually appealing and well-aligned results. This success motivated us to choose depth-conditioned MV-ControlNet as our model. Additional experimental details and results are available in the appendix.
Figure 7: Human preference study on editability and text alignment in five prompt domains of our benchmark. ControlDreamer consistently ranks highest in user preferences across all domains.
Figure 8: The left plot illustrates the 3D model generation results at 45 degrees and 135 degrees under various input conditions. On the right, (a) demonstrates the refinement using Magic3D, and (b) highlights the results achieved with our ControlDreamer.
Figure 9: Comparison with previous text-to-3D pipelines. Hulk’s geometry from Fig. 1 is styled with Ironman and Spiderman’s styles respectively. Magic3D and ProlificDreamer often show texture artifacts due to misalignment between style prompts and the initial geometry. Fantasia3D and MVDream tend to result in oversaturation or residual appearance traits.
4.2 Qualitative Comparison of 3D Generation
We compare against several baselines, namely Magic3D [21], Fantasia3D [5], and ProlificDreamer [42]. All these methods, including our ControlDreamer, refine textured meshes, employing the DMTet [38] algorithm for mesh generation.
The results of our stylization experiments using these baseline models are illustrated in Fig. 9. With Magic3D and ProlificDreamer, we observe some instances of failure, characterized by only partial texture changes in the geometry. This issue arises from a misalignment between the prior knowledge embedded in Stable Diffusion for the style prompt (e.g., the Spiderman is slim) and the geometry from the previous stage. Our MV-ControlNet addresses this misalignment by understanding the depth information of the geometry, thereby enhancing alignment and coherence. While Fantasia3D and MVDream successfully stylize the overall geometry, they tend to produce results with either oversaturation or residual appearance traits from the original geometry prompt, such as persistent green skin color.
Table 2: Comparison of directional CLIP similarity on five different categories.
Domain Animals Characters Foods General Objects
Magic3D 0.218 0.280 0.236 0.247 0.217
Fantasia3D 0.191 0.250 0.256 0.240 0.208
ProlificDreamer 0.224 0.265 0.239 0.231 0.163
MVDream 2-stage 0.160 0.235 0.150 0.211 0.122
ControlDreamer 0.241 0.353 0.270 0.341 0.255
4.3 Quantitative Comparison of 3D Generation
4.3.1 Directional CLIP Similarity
To quantitatively compare the alignment between prompt modifications and image changes in ControlDreamer and MVDream, we evaluate the directional CLIP similarity [9, 16] between first and second stage 3D models on our benchmark dataset. This score measures how well the editing directions of rendered images align with the text directions within the CLIP space. We use texts with three view-oriented postfixes: ’a front view’, ’a side view’, and ’a back view’, utilizing four images from each camera view (0°, 90°, 180°, 270° azimuth angles) for the generated 3D models.
In Table 2, averaged similarities and standard deviations are presented for each domain, across all six prompt candidates and directions. We employ the OpenCLIP ViT-L/14 model [7], trained on the LAION-2B dataset, which is different from the text encoder used in ControlDreamer. Our method demonstrates superior performance compared to other baselines, excelling in generating text-aligned 3D models across all domains in our benchmark. Detailed explanations of the used prompts are presented in Table. A. 3 in Appendix.
4.3.2 Human Evaluations
We assess preference scores using Amazon Mechanical Turk (AMT) across five domains and present them in Fig. 7. As a result, the proposed method consistently achieves the highest preference scores in all domains, demonstrating its effectiveness. This assessment allows for a comparative ranking of MVDream, Fantasia3D, Magic3D, and ProlificDreamer. We assess each methodology using 30 3D models from each domain, with 200 human evaluators participating. Each evaluator encounters three random samples and repeats the evaluation task three times, resulting in about 20 evaluations per domain. For additional details, the templates for the human evaluation are included in the Appendix.
4.4 Ablation Study
4.4.1 Variants of MV-ControlNet
In Fig. 8, we compare the 3D stylization performance of MV-ControlNet variants trained under canny edge, normal, and depth conditions. With canny edge inputs, significant artifacts appear in both RGB and normal maps. When using normal maps, the RGB colors in the 3D models are less photorealistic. However, depth maps consistently produce 3D models without artifacts in both color and normal maps, showcasing their superiority.
4.4.2 MV-ControlNet as a plug-in for prior methods
Additionally, we demonstrate a scenario where MV-ControlNet is used in the refinement stage of existing methodologies, with the identical geometry and style prompts. DreamFusion [29] is employed to create the geometry model. We then compare the performance of MV-ControlNet against Magic3D [21]. As shown in Fig. 8 (b), our approach enhances 3D refinement and effectively avoids semantic drift, highlighting MV-ControlNet’s adaptability as a plug-in component across various methods.
5 Conclusions
We introduce ControlDreamer, a novel pipeline integrating the MV-ControlNet, our novel depth-aware multi-view diffusion model, to enable efficient, text-guided creation of stylized 3D models. This integration effectively generate stylized 3D models by aligning diverse geometries and styles given by textual descriptions. Additionally, our extensive benchmark for 3D style editing, covering a broad range of subjects, establishes a new standard in the field. Comparative analysis confirms that ControlDreamer outperforms existing text-to-3D methods in stylizing 3D models.
6 Limitations and Future Works
While ControlDreamer excels in generating stylized 3D models, it faces several limitations. Errors in pre-trained depth estimators can cause hole artifacts from the geometry stage to persist in the style stage. Another constraint is our experiment’s 256x256 resolution limit, due to MVDream’s training at this resolution. Integrating models like SDXL [28], trained on 1024x1024 images, could be a solution for higher-resolution outcomes. Future work will likely focus on developing an image-conditioned MV-ControlNet and exploring data scaling strategies.
Appendix 0.A Additional Experiments
0.A.1 Qualitative results
In Fig. 10, we visualize more diverse qualitative results used for evaluating directional CLIP similarities and human evaluations.
0.A.2 Ablation Studies
0.A.2.1 Exploring 3D Representations in Style Stages
Fig. 11 illustrates the approach of refining 3D models with NeRF instead of DMTet during the style stage. We observe that volume rendering with NeRF often leads to substantial alterations in the overall geometry, especially when geometry and style texts are semantically different, resulting in instability in depth-aware score distillation. Conversely, employing DMTet, coupled with our timestep annealing approach, proves more effective. It ensures appropriate alterations in the overall texture while maintaining the geometry aligned with the source mesh, leading to more seamless style generation. For NeRF-based 3D stylization, we sample diffusion timesteps from a range of [0.02, 0.5] to minimize significant geometry changes. MV-ControlNet often ignores camera conditions, leading to the creation of multiple artifacts, which further supports the use of DMTet in the style stage.
0.A.2.2 Advantages of MV-ControlNet in multi-view image generation
Fig. 12 illustrates the advantages of the proposed MV-ControlNet through 2D image generation. In Fig. 12 (A), conventional methods exhibit significant semantic variations depending on random seed and camera variations, highly deviating from the original identity of the source image. In contrast, our proposed method demonstrates a stronger alignment with the depth condition, ensuring consistent object-centric generation quality even with different camera parameters and random seeds.
In Fig. 12 (B), when providing text with complex concepts, as shown in (b), the baseline method, likewise as we explained in Fig. 1, displays stereotypical and geometry-biased generation which hampers the image generation quality. On the other hand, our method consistently generates high-quality images in line with the given text without forgetting important identities from the source (e.g., the backpack of an astronaut).
In summary, the proposed MV-ControlNet is more strongly grounded to depth conditions and adept at reflecting complex concepts of the text for the generation of multi-view images.
0.A.3 Effectiveness of Prompt Engineering for 3D Generation
When we use user-defined text prompts not included in the original training set of the MVDream model, we observe artifact generation in the learned source geometry, as shown in Fig. 13. However, we find that careful prompt engineering helps mitigate these artifacts. An example of such a refined text in our experiment is ‘Hulk, muscular, green, 4K, photorealistic’.
0.A.4 Geometry bias in Objaverse dataset
We include additional results on the geometry bias discussed in the introduction of the main paper. Fig. 14 showcases text-based 3D model retrieval on Objaverse dataset, providing a straightforward visualization of the frequency of certain geometries. Notably, we show in the main text that 3D models predominantly associated with a shield are frequently generated from the prompt ‘Captain America’.
0.A.5 Enhanced Stylization in the Image-to-3D Framework
Our methodology exhibits exceptional stylization capabilities, independent of the initialized geometry. Building on this strength, we extend our experiments to an image-to-3D framework [41] to evaluate its editing capabilities. On 3D models generated by ImageDream  [41], we modify the style with our MV-ControlNet. The results, as depicted in Fig. 15, highlight the adaptability of our approach to the image-to-3D conversion process.
0.A.6 Multi-view image editing
In Fig. 16, we demonstrate the editing results of multi-view images using various prompts. For instance, with ’Superman’ as the geometry prompt, we add a cloak to the initial geometry. While the Prompt-to-Prompt (P2P) approach [12] relies on cross-attention control, our method incorporates depth conditions as an additional input to better align with the target style. P2P struggles to produce photorealistic images for counterfactual concepts like Spiderman wearing a cloak. In contrast, our method consistently achieves high-quality results with faithful text-alignments.
0.A.7 Failure cases
In Fig. 17, we present examples of our failure cases. We empirically observe that when the style prompt significantly differs from the initial geometry, our method can struggle to generate visually appealing 3D models. This limitation likely stems from the insufficient capability of the pre-trained diffusion model, namely MVDream, to compose two distinct concepts.
Appendix 0.B Experiment Details
0.B.1 Baselines
We utilize threestudio [10] implementation of Magic3D [21], Fantasia3D [5], and ProlificDreamer [42] as our baselines. Similar to ControlDreamer, all baselines undergo a refinement process on textured meshes derived from the pre-trained MVDream [39] using the DMTet [38] algorithm. Fantasia3D comprises two stages: geometry modeling and appearance modeling, and we use appearance modeling to modify the textured mesh. In the case of the ProlificDreamer, we employ a textured mesh fine-tuning stage for stylization. Additionally, despite ProlificDreamer not utilizing a separate shading model, we apply shading to ensure fair comparisons with other baselines. Our paired geometry and style prompts are shown in Table 3.
Table 3: Comprehensive descriptions of the utilized geometry and style prompts.
Domain Geometry texts Style texts
Animals A bald eagle carved out of wood A highly-detailed rendering of a magpie, ultra hd, realistic, vivid colors
A high-resolution photo of a frog, 3d asset A high-resolution photo of a toad, 3d asset
A cute shark, plush toy, ultra realistic, 4K, HD A highly-detailed photo of a tuna, 4k, HD
A cute shark, plush toy, ultra realistic, 4K, HD A highly-detailed rendering of a killer whale, 4k, HD
A high-resolution rendering of a Shiba Inu, 3d asset A white tiger, super detailed, best quality, 4K, HD
A high-resolution photo of a British Shorthair, 3d asset A highly-detailed photo of a lioness, 4K, HD
Character A high-resolution rendering of a Hulk, 3d asset A highly-detailed 3d rendering of a Superman
A high-resolution rendering of a Hulk, 3d asset A high-resolution rendering of an Iron Man, 3d asset
A high-resolution rendering of a Hulk, 3d asset A highly-detailed photo of a Spider-Man, 4K, HD
A high-resolution rendering of a Hulk, 3d asset Renaissance sculpture of Michelangelo’s David, Masterpiece
A high-resolution rendering of a Hulk, 3d asset A highly-detailed 3d rendering of Thanos wearing the infinity gauntlet
A high-resolution rendering of a Hulk, 3d asset A highly-detailed 3d rendering of a Captain America
Foods A delicious croissant, realistic, 4K, HD A delicious red velvet cake, realistic, 4K, HD
A delicious croissant, realistic, 4K, HD A delicious pecan pie, realistic, 4K, HD
A delicious croissant, realistic, 4K, HD A delicious oreo cake, realistic, 4K, HD
wedding cake, 4k stack of macarons, 4k
wedding cake, 4k strawberry cake, 4k
wedding cake, 4k cheeseburger, 4k
General a cat riding a scooter like a human panda riding a scooter like a human, 4k
a cat riding a scooter like a human tiger riding a scooter like a human, 4k
a cat riding a scooter like a human corgi riding a scooter like a human, 4k
An astronaut riding a horse A high-resolution rendering of an armored cavalry medieval knight, 3d asset
An astronaut riding a horse Gandalf riding a donkey, fairy tale style, 4K, HD
An astronaut riding a horse A beautiful portrait of a princess riding an unicorn, fantasy, HD
Objects A cute shark, plush toy, ultra realistic, 4K, HD A DSLR photo of a submarine, 4k, HD
Battletech Zeus with a sword!, tabletop, miniature, Optimus Prime fighting, super detailed, best quality, 4K, HD
battletech, miniatures, wargames, 3d asset
Battletech Zeus with a sword!, tabletop, miniature, A high-resolution rendering of a Gundam, 3d asset
battletech, miniatures, wargames, 3d asset
Nike Jordan shoe, 4k leather shoe, 4k
Nike Jordan shoe, 4k shoe with flower printed on it, 4k
3d rendering of a mug of hot chocolate with whipped cream A highly-detailed photo of a mug with Starbucks logo, 4K, HD
0.B.2 Generated images used for MV-ControlNet training
In Fig. 18 and Fig. 19, we present additional examples used in training our MV-ControlNet. Both figures illustrate samples created through random sampling, with the camera views for each sample also randomized. These figures clearly show that samples with filtered $100K$ text corpus include a wide range of 3D assets, such as objects, animals, and characters. We utilized these high-quality multi-view images to train MV-ControlNet and thoroughly assessed its performance using diverse user-crafted prompts.
0.B.3 Human evaluations
Our human evaluation spans paired prompts from five domains, each involving six prompts for five methodologies: Fantasia3D, Magic3D, ProlificDreamer, MVDream, and our approach. Evaluations focus on editability and text alignment, with specific questions outlined in the table below.
Your task is to evaluate the quality of 4.0-second-long food video clips. Each clip showcases the results of editing 3D models using text prompts. Please choose the most favorable video based on the following criteria: 1) Overall editing quantity: Has the edited video completely changed in visual appearance compared to the Source? 2) Textual alignment: How well does the text description align with the edited video?
Figure 10: Additional qualitative results. We generate multiple styles on top of various source 3D models.
Figure 11: We visualize the results for both the two-stage MVDream and our ControlDreamer method in the style stage using NeRF representation. Notably, with NeRF as the representation, the geometry undergoes significant changes. This substantial alteration in geometry and depth information during NeRF training leads to instability in our depth-aware score distillation process, resulting in the generation of gravel-like artifacts in our results.
Figure 12: (A) The first column shows source multi-view images along with their depths. We showcase a comparison between images generated by MVDream (second column) and MV-ControlNet (third column), each originating from two different initial noises. MV-ControlNet’s depth-aware generation consistently preserves the source’s geometry, effectively handling variations in initial noise. (B) Our model excels in generating unique images, such as a knight with a backpack, and manages counterfactual prompts like ‘climber zebra’. It produces high-quality images, outperforming MVDream, which struggles with such tasks.
Figure 13: The comparison shows 3D models generated from user-provided texts versus ChatGPT-recommended ones. The left column has models from user texts; the right, those recommended by ChatGPT. Notably, user texts may cause artifacts, as seen in normal maps, whereas ChatGPT’s refined texts result in higher-quality 3D models.
Figure 14: Intriguingly, a search for ‘Captain America’ on https://objaverse.allenai.org/explore primarily yields content related to shields.
Figure 15: Visualization of our stylization process within the image-to-3D pipeline. The source image prompts chosen for this demonstration were ‘pig’ and ‘crab’. Across these distinct subjects, our method maintained robust editing performance, showcasing its consistency in 3D stylization.
Figure 16: The comparative results on multi-view image editings. Our method consistently generates more photorealistic results than P2P [12].
Figure 17: We visualize failure cases of ControlDreamer. In these examples, we use the Hulk geometry from Fig. 1 of the main manuscript. We observe that when the stylization text significantly deviates from the source geometry, it results in the generation of visually unappealing 3D models.
Figure 18: Examples from our generated dataset used for training MV-ControlNet, utilizing the refined text corpus. Each image was created employing a randomized camera setup.
Figure 19: Additional examples of our generated dataset. A randomized camera is used for the generation of each image.